OpenStudy (peterg2001):

The circumference of a circle varies directly as the diameter of a circle. When the diameter of a circle is 5, the circumference of the circle is approximately 15.7. What is the approximate circumference of a circle with a diameter of 9? Round answer to nearest tenth. 2.9 8.7 28.3 706.5

OpenStudy (peterg2001):

@kevinngo

OpenStudy (mimi_x3):

Well we know the circumference of a circle is 2*pi*r where r=radius the radius is half the diameter So If the diameter is 9 then the radius is 9/2=4.5 So then circumference of the circle is 2*pi*4.5 or in other words pi*9 = 28.27 which can be rounded to 28.3

OpenStudy (mimi_x3):

we can use proportions if we would like since we have another circle to compare to